OpenStudy (anonymous):

solve for x, x^4 -2x^2+1=0

OpenStudy (fibonaccichick666):

can you factor it?

OpenStudy (anonymous):

i dont know where to start :(

OpenStudy (fibonaccichick666):

so it is of the form \[(a+b)(c+d)=0\]

OpenStudy (anonymous):

yer i can do that i think

OpenStudy (fibonaccichick666):

ok so what will be your a?

OpenStudy (anonymous):

x?

OpenStudy (fibonaccichick666):

have you heard of "completing the square"?

OpenStudy (anonymous):

that the worst thing ever but yes

OpenStudy (fibonaccichick666):

that's what you have to do for this problem so to start:

OpenStudy (anonymous):

ok thank you

OpenStudy (fibonaccichick666):

what number multiplies to 1 but adds to -2?

OpenStudy (anonymous):

u tell me

OpenStudy (fibonaccichick666):

nope tell me

OpenStudy (anonymous):

would it be like -1?

OpenStudy (fibonaccichick666):

please plug it in and tell me if it satisfies the two conditions I gave you

OpenStudy (anonymous):

look i got no idea

OpenStudy (fibonaccichick666):

does -1 multiply to 1

OpenStudy (anonymous):

i got it!!!! x=-1 or x=1

OpenStudy (fibonaccichick666):

close, can you multiply those together and tell me if they =1?

OpenStudy (anonymous):

no wouldnt it equal -1?

OpenStudy (fibonaccichick666):

that is correct. and we know from our original eq. that they need to =1. so what two numbers multiply to positive one?

OpenStudy (anonymous):

two negitive = a positive

OpenStudy (fibonaccichick666):

right, so what multiplies to one?

OpenStudy (anonymous):

1

OpenStudy (fibonaccichick666):

but we wanted negative numbers right? because then they need to add to -2

OpenStudy (anonymous):

yer so the second would be -1?

OpenStudy (fibonaccichick666):

but they both need to be negative right?

OpenStudy (anonymous):

right

OpenStudy (fibonaccichick666):

so what two negative numbers multiply to 1 and add to -2?

OpenStudy (anonymous):

-1

OpenStudy (fibonaccichick666):

good! now if you know that (a-1)(b-1)=x^4-2x^2+1=0 what are your a and b?

OpenStudy (fibonaccichick666):

hint: a might equal b

OpenStudy (dwelsher):

1.factor it first 2.(x^2-1)(x^2-1)=====(x-1)(x+1)(x-1)(x+1)=0 3.x=1or-1??????